THIS PROBLEM MUST BE SOLVED USING SYSTEMS.
The manager of a candy shop sells chocolate covered peanuts for $7 per pound and chocolate covered cashews for $14 per pound. The manager wishes to make a 245-pound cashew-peanut mixture that will sell for $9 per pound. How many pounds of each should be used?

Answer by mananth(16946)

x= Chocolate covered peanuts pounds
y= Chocolate covered cashewnuts pounds
Total nuts sold
1 x -1 y = 245 .............1
Cost of nuts
7 x + 14 y = 9*245 .............2
Eliminate y
multiply (1)by 14
Multiply (2) by 1
14 x -14 y = 3430
7 x + 14 y = 2205
Add the two equations
21 x = 5635
/ 21
x = 268
plug value of x in (1)
1 x -1 y = 245
268 -1 y = 245
-1 y = 245 -268
-1 y = -23
y = 23
268 pounds Chocolate covered peanuts
23 pounds Chocolate covered cashewnuts
m.ananth@hotmail.ca